Linux PCMCIA HOWTO

David Hinds, dhinds@allegro.stanford.edu v1.87, 1996/07/29 01:59:46

This document describes how to install and use PCMCIA Card Services for Linux, and answers some frequently asked questions. The latest version of this document can always be found at hyper.stanford.edu

in /pub/pcmcia/doc. An HTML version is at http://hyper.stanford.edu/~dhinds/pcmcia/pcmcia.html

Contents

1 General information and hardware requirements 2 1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Copyright notice and disclaimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 What is the latest version, and where can I get it? . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 What systems are supported? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.5 What PCMCIA cards are supported? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.6 When will card X be supported? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.7 Mailing list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Compilation, installation, and configuration 5 2.1 Prerequisites and kernel setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Site-specific configuration options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4 Problems loading kernel modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5 Problems with the card status change interrupt . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.6 Memory window configuration problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.7 What about installation on Red Hat and Caldera? . . . . . . . . . . . . . . . . . . . . . . . . 10 2.8 Why don’t you distribute PCMCIA binaries? . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.9 Why is the PCMCIA package so darned big? . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3 Usage and features 11 3.1 How do I tell if it is working? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2 Overview of the PCMCIA configuration scripts . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.3 How do I use my PCMCIA network card? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.4 How do I use my PCMCIA modem card? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.5 How do I use my PCMCIA SCSI card? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.6 How do I use my PCMCIA memory card? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.7 How do I use my ATA/IDE card drive? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.8 How do I tell cardmgr how to identify a new card? . . . . . . . . . . . . . . . . . . . . . . . 19 3.9 How do I control which interrupts and ports are used by a device? . . . . . . . . . . . . . . . 19 3.10 When is it safe to insert or eject a PCMCIA card? . . . . . . . . . . . . . . . . . . . . . . . . 20 3.11 How do I unload PCMCIA drivers? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.12 How does Card Services deal with suspend/resume? . . . . . . . . . . . . . . . . . . . . . . . 20 3.13 How do I turn off a PCMCIA card without ejecting it? . . . . . . . . . . . . . . . . . . . . . . 21 3.14 How can I have separate device configurations for home and work? . . . . . . . . . . . . . . . 21

4 Problems with specific cards 21 4.1 Why doesn’t my modem work? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.2 Why doesn’t my network adapter card work? . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.3 How do I select the transceiver type for a network card? . . . . . . . . . . . . . . . . . . . . . 23 4.4 How do I add support for an NE2000-compatible ethernet card? . . . . . . . . . . . . . . . . 24 4.5 Multifunction cards such as the 3Com 3c562 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.6 How do I use my PCMCIA floppy interface? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.7 What’s up with support for Xircom cards? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

5 Debugging tips and programming information 25 5.1 How can I submit a helpful bug report? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.2 Low level PCMCIA debugging aids . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 5.3 How do I write a Card Services driver for card X? . . . . . . . . . . . . . . . . . . . . . . . . 26

1 General information and hardware requirements

1.1 Introduction

Card Services for Linux is a complete PCMCIA support package. It includes a set of loadable kernel modules that implement a version of the PCMCIA Card Services applications program interface, a set of client drivers for specific cards, and a card manager daemon that can respond to card insertion and removal events, loading and unloading drivers on demand. It supports “hot swapping” of PCMCIA cards, so cards can be inserted and ejected at any time.

This software is still under development. It probably contains bugs, and should be used with caution. I’ll do my best to fix problems that are reported to me, but if you don’t tell me, I may never know. If you use this code, I hope you will send me your experiences, good or bad!

If you have any suggestions for how this document could be improved, please let me know

(dhinds@allegro.stanford.edu).

1.2 Copyright notice and disclaimer

Copyright (c) 1996 David A. Hinds

This document may be reproduced or distributed in any form without my prior permission. Modified versions of this document, including translations into other languages, may be freely distributed, provided that they are clearly identified as such, and this copyright is included intact.

This document may be included in commercial distributions without my prior consent. While it is not required, I would like to be informed of such usage. If you intend to incorporate this document in a published work, please contact me to make sure you have the latest available version.

This document is provided “as is”, with no explicit or implied warranties. Use the information in this document at your own risk.

1.3 What is the latest version, and where can I get it?

The current major release of Card Services is version 2.8, and minor updates or bug fixes are numbered 2.8.1, 2.8.2, and so on.

Source code for the latest version is available from hyper.stanford.edu in the /pub/pcmcia directory, as pcmcia-cs-2.8.?.tar.gz. There will usually be several versions here. I generally only keep the latest minor release for a given major release. New major releases may contain relatively untested code, so I also keep the latest version of the previous major release as a relatively stable fallback; the current fallback is 2.7.6. It is up to you to decide which version is more appropriate, but the CHANGES file will summarize the most important differences.

hyper.stanford.edu is mirrored at sunsite.unc.edu in /pub/Linux/kernel/pcmcia. I’ll also try to up-
load major releases to tsx-11.mit.edu under /pub/linux/packages/laptops/pcmcia now and then.

If you do not feel up to compiling the PCMCIA drivers from scratch, pre-compiled drivers are included with current releases of most of the major Linux distributions, including Slackware, Red Hat, Caldera, and Yggdrasil, among others.

1.4 What systems are supported?

This code should run on almost any Linux-capable laptop. All common PCMCIA controllers are supported, including Intel, Cirrus, Vadem, VLSI, Ricoh, and Databook chips. Custom controllers used in IBM and Toshiba laptops are also supported. Several people use the package on desktop systems with PCMCIA card adapters.

The Motorola 6AHC05GA controller used in some Hyundai laptops is not supported. The custom PCMCIA controller in the HP Omnibook 600 is also unsupported.

1.5 What PCMCIA cards are supported?

The current release includes drivers for a variety of ethernet cards, a driver for modem and serial port cards, several SCSI adapter drivers, a driver for ATA/IDE drive cards, and memory card drivers that should support most SRAM cards and some flash cards. The SUPPORTED.CARDS file included with each release of Card Services lists all cards that are known to work in at least one actual system.

The likelihood that a card not on the supported list will work depends on the type of card. Essentially all modems should work with the supplied driver. Some network cards may work if they are OEM versions of supported cards. Other types of IO cards (GPS, frame buffers, sound cards, etc) will not work until someone writes the appropriate drivers.

1.6 When will card X be supported?

Unfortunately, they do not pay me to write device drivers, so if you would like to have a driver for your favorite card, you are probably going to have to do at least some of the work. The SUPPORTED.CARDS file mentions some cards for which driver work is currently in progress. I will try to help where I can.

1.7 Mailing list

I maintain a database and mailing list of Linux PCMCIA users. This is used to announce new releases of the PCMCIA package. To be included, send me the following:

• Your name and email address

• What kind of laptop are you using?

• What PCMCIA controller is reported by the probe command?

• What PCMCIA cards are you using?

• Any special settings you use: compilation options, irq and port settings, /etc/pcmcia/config.opts changes, insmod options, etc.

You can also register via the WWW: see http://hyper.stanford.edu/~dhinds/pcmcia/pcmcia.html for

instructions.

There is also a Linux mailing list devoted to laptop issues, the “linux-laptop” list. For more information, send a message containing the word “help” to majordomo@vger.rutgers.edu.

2 Compilation, installation, and configuration

2.1 Prerequisites and kernel setup

The following things should be installed on your system before you start installing PCMCIA:

• One of the following kernels: 1.2.8 through 1.2.13, 1.3.30, 1.3.37, 1.3.39 through 1.3.99, 1.99.* (i.e., pre-2.0), or 2.0.*.

• A current set of module utilities.

• (Optional) the “Forms” X11 user interface toolkit.

The latest version requires a kernel version 1.2.8 or higher, or a development kernel 1.3.30 or higher. 1.3.38 is definitely broken, and 1.3.31 to 1.3.36 are untested. It also requires a relatively recent set of module utilities. There are no kernel patches specifically for PCMCIA.

You need to have a complete linux source tree for your kernel, not just an up-to-date kernel image, to compile the PCMCIA package. The PCMCIA modules contain some references to kernel source files. While you may want to build a new kernel to remove unnecessary drivers, installing PCMCIA does not require you to do so.

Current “stable” kernel sources and patches are available from sunsite.unc.edu in /pub/Linux/kernel/v1.2, or from

tsx-11.mit.edu in /pub/linux/sources/system/v1.2. Current module utilities can be found in the same

place, in modules-1.2.8.tgz. Development kernels can be found in the corresponding v1.3 subdirectories. Some recent development kernels and compiler releases interact badly with older sets of module tools. The latest versions of the module utilities can be found at <http://www.pi.se/blox/modules>.

When configuring your kernel, if you plan on using a PCMCIA ethernet card, you should turn on networking support but turn off the normal Linux network card drivers, including the “pocket and portable adapters”. The PCMCIA network card drivers are all implemented as loadable modules. Any drivers compiled into your kernel will only waste space.

If you want to use SLIP, PPP, or PLIP, you do need to either configure your kernel with these enabled, or use the loadable module versions of these drivers. There is an unfortunate deficiency in the kernel config process in 1.2.X kernels, in that it is not possible to set configuration options (like SLIP compression) for a loadable module, so it is probably better to just link SLIP into the kernel if you need it.

If you want to use a PCMCIA token ring adapter, your kernel needs to be configured with “Token Ring driver support” (CONFIG TR) enabled, though you should leave CONFIG IBMTR off.

If you will be using a PCMCIA SCSI adapter, you should enable CONFIG SCSI when configuring your kernel. Also, enable any top level drivers (SCSI disk, tape, cdrom, generic) that you expect to use. All low-level drivers for particular host adapters should be disabled, as they will just take up space.

If you want to modularize a driver that is needed for a PCMCIA device, you must modify /etc/pcmcia/config to specify what modules need to be loaded for what card types. For example, if the serial driver is modularized, then you could change the serial device definition to:

     device "serial_cs"

       class "serial" module "serial", "serial_cs"

If your kernel is compiled with CONFIG MODVERSIONS enabled, for kernel symbol version checking, the configure script will check for the existence of /usr/include/linux/modversions.h, the symbol version database. This is created by running “make dep” in the kernel source tree.

This package includes an X-based card status utility called cardinfo. This utility is based on a public domain user interface toolkit called the Forms Library, which you will need to install before building cardinfo. A binary distribution is on hyper.stanford.edu in /pub/pcmcia/extras: there are both a.out and ELF versions of the library.

2.2 Installation

Here is a synopsis of the installation process:

• Unpack pcmcia-cs-2.8.?.tar.gz in /usr/src.

   • Run “make config” in the new pcmcia-cs-2.8.? directory.
   • Run “make all”, then “make install”.

• Customize /etc/rc.d/rc.pcmcia and the option files in /etc/pcmcia for your site.

Running “make config” prompts for a few configuration options, and checks out your system to verify that it satisfies all prerequisites for installing PCMCIA support. In most cases, you’ll be able to just accept all the default configuration options. Be sure to carefully check the output of this command in case there are problems.

If you are compiling the PCMCIA stuff for installation on another machine, specify an alternate target directory when prompted by the configure script. This should be an absolute path. All the PCMCIA tools will be installed relative to this directory. You will then be able to tar this directory tree and copy to your target machine, and unpack relative to its root directory to install everything in the proper places.

If you are cross compiling on another machine, you may want to specify alternate names for the compiler and linker. This may also be helpful on mixed a.out and ELF systems. The script will also prompt for additional compiler flags for debugging.

There are a few kernel configuration options that affect the PCMCIA tools. The configuration script can deduce these from the running kernel (the most common case). Alternatively, if you are compiling for installation on another machine, it can read the configuration from a kernel source tree, or each option can be set interactively.

Running “make all” followed by “make install” will build and then install the kernel modules and utility programs. Kernel modules are installed under /lib/modules/<version>/pcmcia. The cardmgr and cardctl programs are installed in /sbin. If cardinfo is built, it is installed in /usr/bin/X11.

Configuration files will be installed in the /etc/pcmcia directory. If you are installing over an older version, your old config scripts will be backed up before being replaced. The saved scripts will be given extensions like *.~1~, *.~2~, and so on.

If you don’t know what kind of PCMCIA controller chip you have, you can use the probe utility in the cardmgr/ subdirectory to determine this. There are two major types: the Databook TCIC-2 type and the Intel i82365SL-compatible type.

A user-level daemon processes card insertion and removal events. This is called cardmgr. It is similar in function to Barry Jaspan’s pcmciad in earlier PCMCIA releases. Cardmgr reads a configuration file describing known PCMCIA cards from /etc/pcmcia/config. This file also specifies what resources can be allocated for use by PCMCIA devices, and may need to be customized for your system. See the pcmcia man page for more information about this file.

The script rc.pcmcia, installed in /etc/rc.d, controls starting up and shutting down the PCMCIA system. “make install” will use the probe command to determine your controller type and modify rc.pcmcia appropriately. You should add a line to your system startup file /etc/rc.d/rc.M to invoke this:

     /etc/rc.d/rc.pcmcia start

In a few cases, the probe command will be unable to determine your controller type automatically. The Tadpole P1000 and some other PCI-based laptops have a special Cirrus PCI-to-PCMCIA bridge chip that can’t be detected by probe. If you have one of these systems, you’ll need to edit rc.pcmcia by hand to load the i82365 module. If you have a Halikan NBD 486 system, it has a TCIC-2 controller at an unusual location: you’ll need to edit rc.pcmcia to load the tcic module, as well as setting the PCIC OPTS parameter

to “tcic base=0x02c0”.

2.3 Site-specific configuration options

Card Services should automatically avoid allocating IO ports and interrupts already in use by other standard devices. It will also attempt to detect conflicts with unknown devices, but this is not completely reliable. In some cases, you may need to explicitly exclude resources for a device in /etc/pcmcia/config.opts.

Here are some resource settings for specific laptop types.

• On the AMS SoundPro, exclude irq 10.

• On some AMS TravelPro 5300 models, use memory 0xc8000-0xcffff.

• On the BMX 486DX2-66, exclude irq 5, irq 9.

• On the Chicony NB5, use memory 0xda000-0xdffff.

• On the HP Omnibook 4000C, exclude port 0x300-0x30f.

• On the NEC Versa M, exclude irq 9, port 0x2e0-2ff.

• On the NEC Versa P/75, exclude irq 5, irq 9.

• On the NEC Versa S, exclude irq 9, irq 12.

• On the ProStar 9200, Altima Virage, and Acquiline Hurricane DX4-100, exclude irq 5, port 0x330-0x35f. Maybe use memory 0xd8000-0xdffff.

• On the TI TravelMate 5000, use memory 0xd4000-0xdffff.

• On the Toshiba T4900 CT, exclude irq 5, port 0x2e0-0x2e8, port 0x330-0x338.

• On the Twinhead 5100, HP 4000, Sharp PC-8700 and PC-8900, exclude irq 9 (sound), irq 12.

• On an MPC 800 Series, exclude irq 5, port 0x300-0x30f for the CD-ROM.

Some PCMCIA controllers have optional features that may or may not be implemented in a particular system. It is generally impossible for a socket driver to detect if these features are implemented. Check the man page for your driver to see what optional features may be enabled.

The low level socket drivers, tcic and i82365, have numerous bus timing parameters that may need to be adjusted for systems with particularly fast processors. Symptoms of timing problems include card recognition problems, lock-ups under heavy loads, high error rates, or poor device performance. Check the corresponding man pages for more details, but here is a brief summary:

• Cirrus controllers have numerous configurable timing parameters. The most important is the freq bypass flag which changes the multiplier for the PCMCIA bus clock to slow down all operations.

• The Cirrus PD6729 PCI controller has the fast pci flag, which should be set if the PCI bus speed is greater than 25 MHz.

• For Vadem VG-468 controllers and Databook TCIC-2 controllers, the async clock flag changes the relative clocking of PCMCIA bus and host bus cycles. Setting this flag adds extra wait states to some operations. However, I have yet to hear of a laptop that needs this.

• The pcmcia core module has the cis speed parameter for changing the memory speed used for accessing a card’s Card Information Structure (CIS). On some systems with fast bus clocks, increasing this parameter (i.e., slowing down card accesses) may be beneficial.

• This isn’t a timing issue, but if you have more than one PCMCIA controller in your system or extra sockets in a docking station, the i82365 module should be loaded with the extra sockets parameter set to 1.

All these options should be configured by modifying the top of /etc/rc.d/rc.pcmcia. For example:

     # Should be either i82365 or tcic
     PCIC=i82365
     # Put socket driver timing parameters here
     PCIC_OPTS="cmd_time=12"
     # Put pcmcia_core options here
     CORE_OPTS="cis_speed=500"

Here are some timing settings for specific systems:

• On the ARM Pentium-90 or Midwest Micro Soundbook Plus, use “freq bypass=1 cmd time=8”.

• On a Midwest Micro Soundbook Elite, use “cmd time=12”.

• On a Gateway Liberty, try “cmd time=16”.

On some systems using Cirrus controllers, including the NEC Versa M, the BIOS puts the controller in a special suspended state at system startup time. On these systems, the probe command will fail to find any known PCMCIA controller. If this happens, edit /etc/rc.d/rc.pcmcia by hand as follows:

     # Should be either i82365 or tcic
     PCIC=i82365
     # Put socket driver timing parameters here
     PCIC_OPTS="wakeup=1"

2.4 Problems loading kernel modules

The configure script will normally ensure that the PCMCIA modules are compatible with your kernel. So, module loading problems generally indicate that the user has interfered with the normal installation process in some way. Some module loading errors are sent directly to the Linux console. Other errors are recorded in the system log file, normally /usr/adm/messages or /var/log/messages. To track down a problem, be sure to check both locations, to pin down which module is actually causing trouble.

Some of the PCMCIA modules require kernel services that may or may not be present, depending on kernel configuration. For instance, the SCSI card drivers require that the kernel be configured with SCSI support, and the network drivers require a networking kernel. If a kernel lacks a necessary feature, insmod may report undefined symbols and refuse to load a module.

If insmod reports “wrong version” errors, it means that the module was compiled for a different kernel version than your system is actually running. This might occur if modules compiled on one machine are copied to another machine with a different configuration, or if the kernel is reconfigured after PCMCIA is installed.

Another source of module loading errors is when the modules and kernel were compiled with different settings of CONFIG MODVERSIONS. If a module with version checking is loaded against a kernel without version checking, insmod will complain about undefined symbols.

Finally, relatively recent binutils releases are incompatible with older versions of the module utilities, and can cause module version incompatibilities to be reported. The most common symptom is complaints about “gcc compiled” being undefined. If you get these errors, upgrade to the latest module utilities, available

from <http://www.pi.se/blox/modules>.

2.5 Problems with the card status change interrupt

In most cases, the socket driver (i82365 or tcic) will automatically probe and select an appropriate interrupt to signal card status changes. The automatic interrupt probe doesn’t work on some Intel-compatible controllers, including Cirrus chips and the chips used in some IBM ThinkPads. If a device is inactive at probe time, its interrupt may also appear to be available. In these cases, the socket driver may pick an interrupt that is used by another device.

With the i82365 and tcic drivers, the irq mask option can be used to limit the interrupts that will be tested. This mask limits the set of interrupts that can be used by PCMCIA cards as well as for monitoring card status changes. The cs irq option can also be used to explicitly set the interrupt to be used for monitoring card status changes.

If you can’t find an interrupt number that works, there is also a polled status mode: both i82365 and tcic will accept a poll interval=100 option, to poll for card status changes once per second. This option should also be used if your system has a shortage of interrupts available for use by PCMCIA cards. Especially for systems with more than one PCMCIA controller, there is little point in dedicating interrupts for monitoring card status changes.

All these options should be set in the PCIC OPTS= line in either /etc/rc.d/rc.pcmcia or /etc/sysconfig/pcmcia, depending on your site setup.

The most common problem of this type seems to be a conflict with a PS/2 pointer device on interrupt 12. In this case, edit rc.pcmcia so the PCIC OPTS= line reads:

     PCIC_OPTS="irq_mask=0xefff"

2.6 Memory window configuration problems

By default, the PCMCIA drivers use the address region 0xd0000-0xdffff to access PCMCIA devices. This memory window is specified in /etc/pcmcia/config.opts. If this address region is used by other devices in your system, cards may not be identified correctly. With chipsets that support it, conflicts can also result from memory shadowing in this region. If you find that all your cards are always misidentified as memory cards, verify that shadowing is disabled in your system’s hardware setup. As there is no reliable way to automatically identify memory window conflicts, finding a good window may require some experimentation. Some alternative windows to try are 0xd8000-0xdffff, 0xc0000-0xcffff, and 0xc8000-0xcffff.

If you have DOS PCMCIA drivers, you may be able to deduce what memory region those drivers use. Note that DOS memory addresses are often specified in “segment” form, which leaves off the final hex digit (so an absolute address of 0xd0000 would be given as 0xd000). Be sure to add the extra digit back when making

changes to /etc/pcmcia/config.opts.

2.7 What about installation on Red Hat and Caldera?

Red Hat and Caldera have a System V-ish arrangement for system startup files. The PCMCIA installation scripts will automatically detect this and adjust accordingly. The rc.pcmcia script will be installed as /etc/rc.d/init.d/pcmcia. There is no need to edit any of the Caldera startup scripts to enable PCMCIA: it will happen automatically.

A separate configuration file, /etc/sysconfig/pcmcia, will be created for startup options. If you need to change any module options (like the PCIC= or PCIC OPTS= settings), edit this config file rather than the actual PCMCIA startup script. This file will not be overwritten by subsequent installs.

Previous releases used the /etc/sysconfig/pcmcia-scripts directory in place of /etc/pcmcia on these platforms. The current release instead uses /etc/pcmcia for all systems, and will move an existing

/etc/sysconfig/pcmcia-scripts to /etc/pcmcia.

2.8 Why don’t you distribute PCMCIA binaries?

For me, distributing binaries is a significant hassle. It is complicated because some features can only be selected at compile time, and because the PCMCIA modules are somewhat dependent on having the “right” kernel configuration. So, I would probably need to distribute precompiled modules along with matching kernels. Beyond this, the greatest need for precompiled modules is when installing Linux on a clean system. This typically requires setting up PCMCIA so that it can be used in the installation process for a particular Linux distribution. Each Linux distribution has its own procedures, and it is not feasible for me to provide boot and root disks for even just the common combinations of drivers and distributions.

PCMCIA is now a part of many of the major Linux distributions, including Red Hat, Caldera, Slackware, Yggdrasil, Craftworks, and Nascent Technology.

2.9 Why is the PCMCIA package so darned big?

Well, first of all, it isn’t actually that large. All the driver modules together take up about 200K of disk space. The utility programs add up to about 70K, and the stuff in /etc/pcmcia is about 30K. When running, the core PCMCIA modules take up 48K of system memory. The cardmgr daemon will generally be swapped out except when cards are inserted or removed. The total package size is not much different from DOS Card Services implementations.

Compared to DOS “point enablers”, this may still seem like a lot of overhead, especially for people that don’t plan on using many of the features of PCMCIA, such as power management or hot swapping. Point enablers can be tiny because they generally support only one or a small set of cards, and also generally support a restricted set of PCMCIA controllers. If someone were to write a genuinely “generic” modem enabler, it would end up incorporating much of the functionality of Card Services, to handle cards from different vendors and the full range of PCMCIA controller variants.

3 Usage and features

3.1 How do I tell if it is working?

The cardmgr daemon normally beeps when a card is inserted, and the tone of the beeps indicates the status of the newly inserted card. Two high beeps indicate the card was identified and configured successfully. A high beep followed by a lower beep indicates that the card was identified, but could not be configured for some reason. One low beep indicates that the card could not be identified.

If you are running X, the new cardinfo utility produces a slick graphical display showing the current status of all PCMCIA sockets.

If the modules are all loaded correctly, the output of the lsmod command should look like the following, with no cards inserted:

     Module:        #pages:  Used by:
     ds                 2
     i82365             3
     pcmcia_core        7    [ds i82365]

All the PCMCIA modules and the cardmgr daemon send status messages to the system log. This will usually be /var/log/messages or /usr/adm/messages. This file should be the first place to look when tracking down a problem. When submitting a bug report, always include the contents of this file. Cardmgr also records some current device information for each socket in /var/run/stab. Here is a sample /var/run/stab listing:

     Socket 0: Adaptec APA-1460 SlimSCSI
     0       scsi    aha152x_cs      0       sda     8       0
     0       scsi    aha152x_cs      1       scd0    11      0
     Socket 1: Serial or Modem Card
     1       serial  serial_cs       0       cua1    5       65

For the lines describing devices, the first field is the socket, the second is the device class, the third is the driver name, the fourth is used to number multiple devices associated with the same driver, the fifth is the device name, and the final two fields are the major and minor device numbers for this device.

3.2 Overview of the PCMCIA configuration scripts

Each PCMCIA device has an associated “class” that describes how it should be configured and managed. Classes are associated with device drivers in /etc/pcmcia/config. There are currently five IO device classes (network, SCSI, cdrom, fixed disk, and serial) and three memory device classes (FTL, memory, and pcmem). For each class, there are two scripts in /etc/pcmcia/config: a main configuration script (i.e., /etc/pcmcia/scsi for SCSI devices), and an options script (i.e., /etc/pcmcia/scsi.opts). The main script for a device will be invoked to configure that device when a card is inserted, and to shut down the device when the card is removed. For cards with several associated devices, the script will be invoked for each device.

The config scripts start by extracting some information about a device from /var/run/stab. Each script constructs a “device address”, that uniquely describes the device it has been asked to configure, in the ADDRESS variable. This is passed to the *.opts script, which should return information about how a device at this address should be configured. For some devices, the device address is just the socket number. For others, it includes extra information that may be useful in deciding how to configure the device. For example, network devices pass their hardware ethernet address as part of the device address, so the network.opts script could use this to select from several different configurations.

The first part of all device addresses is the current PCMCIA “scheme”. This parameter is used to support multiple sets of device configurations based on a single external user-specified variable. One use of schemes would be to have a “home” scheme, and a “work” scheme, which would include different sets of network configuration parameters. The current scheme is selected using the cardctl command. The default if no scheme is set is “default”.

As a general rule, when configuring Linux for a laptop, PCMCIA devices should only be configured from the PCMCIA device scripts. Do not try to configure a PCMCIA device the same way you would configure a permanently attached device.

3.3 How do I use my PCMCIA network card?

Linux ethernet-type network interfaces normally have names like eth0, eth1, and so on. Token-ring adapters are handled similarly, however they are named tr0, tr1, and so on. The ifconfig command is used to view or modify the state of a network interface. A peculiarity of Linux is that network interfaces do not have corresponding device files under /dev, so don’t be surprised when you can’t find them.

When a PCMCIA ethernet card is detected, it will be assigned the first free interface name, which will probably be eth0. Cardmgr will run the /etc/pcmcia/network script to configure the interface.

Do not configure your PCMCIA ethernet card in /etc/rc.d/rc.inet1, since the card may not be present when this script is executed. Comment out everything except the loopback stuff in rc.inet1. If your system has an automatic network configuration procedure, you should indicate that you do not have a network card installed. Instead, edit the /etc/pcmcia/network.opts file to match your local network setup. The network and network.opts scripts will be executed only when your ethernet card is actually present.

The device address passed to network.opts consists of three comma-separated fields: the scheme, the socket number, the device instance, and the card’s hardware ethernet address. The device instance is used to number devices for cards that have several network interfaces, so it will usually be 0. If you have several network cards used for different purposes, one option would be to configure the cards based on socket position, as in:

     case "$ADDRESS" in
     *,0,*,*)
         # definitions for network card in socket 0
         ;;
     *,1,*,*)
         # definitions for network card in socket 1
         ;;
     esac

Alternatively, they could be configured using their hardware addresses, as in:

     case "$ADDRESS" in
     *,*,*,00:80:C8:76:00:B1)
         # definitions for a D-Link card
         ;;
     *,*,*,08:00:5A:44:80:01)
         # definitions for an IBM card
     esac

To automatically mount and unmount NFS filesystems, first add all these filesystems to /etc/fstab, but include noauto in the mount options. In network.opts, list the filesystem mount points in the MOUNTS variable. It is especially important to use either cardctl or cardinfo to shut down a network card when NFS mounts are configured this way. It is not possible to cleanly unmount NFS filesystems if a network card is simply ejected without warning.

3.4 How do I use my PCMCIA modem card?

Linux serial devices are accessed via the /dev/cua* and /dev/ttyS* special device files. The ttyS* devices are for incoming connections, such as directly connected terminals. The cua* devices are for outgoing connections, such as modems. The configuration of a serial device can be examined and modified with the

setserial command.

When a PCMCIA serial or modem card is detected, it will be assigned to the first available serial device slot. This will usually be /dev/cua1 or /dev/cua2, depending on the number of built-in serial ports. The default serial device option script, /etc/pcmcia/serial.opts, will link the appropriate device file to /dev/modem as a convenience.

Do not try to use /etc/rc.d/rc.serial to configure a PCMCIA modem. This script should only be used to configure non-removable devices. Modify /etc/pcmcia/serial.opts if you want to do anything special to set up your modem.

The device address passed to serial.opts has three comma-separated fields: the first is the scheme, the second is the socket number, and the third is the device instance. The device instance may take several values for cards that support multiple serial ports, but for single-port cards, it will always be 0. If you commonly use more than one PCMCIA modem, you may want to specify different settings based on socket position, as in:

     case "$ADDRESS" in
     *,0,*)
         # Options for modem in socket 0
         ;;
     *,1,*)
         # Options for modem in socket 1
         ;;
     esac

If a PCMCIA modem is already configured when Linux boots, it may be incorrectly identified as an ordinary built-in serial port. This is harmless, however, when the PCMCIA drivers take control of the modem, it will be assigned a different device slot. It is best to either parse /var/run/stab or use /dev/modem, rather than expecting a PCMCIA modem to always have the same device assignment.

If you configure your kernel to load the basic Linux serial port driver as a module, you must edit /etc/pcmcia/config to indicate that this module must be loaded. Edit the serial device entry to read:

     device "serial_cs"
       class "serial" module "char/serial", "serial_cs"

3.5 How do I use my PCMCIA SCSI card?

All the currently supported PCMCIA SCSI cards are work-alikes of one of the following ISA bus cards: the Qlogic, the Adaptec AHA-152X, or the Future Domain TMC-16x0. The PCMCIA drivers are built by linking some PCMCIA-specific code (in qlogic cs.c, toaster cs.c, or fdomain cs.c) with the normal Linux SCSI driver.

When a new SCSI host adapter is detected, the SCSI drivers will probe for devices. Check the system log to make sure your devices are detected properly. New SCSI devices will be assigned to the first available SCSI device files. The first SCSI disk will be /dev/sda, the first SCSI tape will be /dev/st0, and the first CD-ROM will be /dev/scd0.

With 1.3.X kernels, the PCMCIA core drivers are able to find out from the kernel which SCSI devices are connected to a card. They will be listed in /var/run/stab, and the SCSI configuration script, /etc/pcmcia/scsi, will be called once for each attached device, to either configure or shut down that device. The default script does not take any actions to configure SCSI devices, but will properly unmount filesystems on SCSI devices when a card is removed.

With 1.2.X kernels, the PCMCIA drivers cannot automatically deduce which devices are associated with a particular SCSI adapter. Instead, if you have one normal SCSI device configuration, you may list these devices in /etc/pcmcia/scsi.opts. For example, if you normally have a SCSI disk and a CD-ROM, you would use:

     # For 1.2 kernels: list of attached devices
     SCSI_DEVICES="sda scd0"

The device addresses passed to scsi.opts are complicated, because of the variety of things that can be attached to a SCSI adapter. Addresses consist of either six or seven comma-separated fields: the current scheme, the device type, the socket number, the SCSI channel, ID, and logical unit number, and optionally, the partition number. The device type will be “sd” for disks, “st” for tapes, “sr” for CD-ROM devices, and “sg” for generic SCSI devices. For most setups, the SCSI channel and logical unit number will be 0. For disk devices with several partitions, scsi.opts will first be called for the whole device, with a five-field address. The script should set the PARTS variable to a list of partitions. Then, scsi.opts will be called for each partition, with the longer seven-field addresses. For example, here is a script for configuring a disk device at SCSI ID 3, with two partitions, and a CD-ROM at SCSI ID 6:

     case "$ADDRESS" in
     *,sd,*,0,3,0)
         # This device has two partitions...
         PARTS="1 2"
         ;;
     *,sd,*,0,3,0,1)
         # Options for partition 1:
         #  update /etc/fstab, and mount an ext2 fs on /usr1
         DO_FSTAB="y" ; DO_FSCK="y" ; DO_MOUNT="y"
         FSTYPE="ext2"
         OPTS=""

         MOUNTPT="/usr1"
         ;;
     *,sd,*,0,3,0,2)
         # Options for partition 2:
         #  update /etc/fstab, and mount an MS-DOS fs on /usr2
         DO_FSTAB="y" ; DO_FSCK="y" ; DO_MOUNT="y"
         FSTYPE="msdos"
         OPTS=""
         MOUNTPT="/usr2"
         ;;
     *,sr,*,0,6,0)
         # Options for CD-ROM at SCSI ID 6
         PARTS=""
         DO_FSTAB="y" ; DO_FSCK="n" ; DO_MOUNT="y"
         FSTYPE="iso9660"
         OPTS="ro"
         MOUNTPT="/cdrom"
         ;;
     esac

If your kernel does not have a top-level driver (disk, tape, etc) for a particular SCSI device, then the device will not be configured by the PCMCIA drivers. As a side effect, the device’s name in /var/run/stab will be something like “sd#nnnn” where “nnnn” is a four-digit hex number. This happens when cardmgr is unable to translate a SCSI device ID into a corresponding Linux device name.

It is possible to modularize the top-level SCSI drivers so that they are only loaded when a PCMCIA SCSI adapter is detected. To do so, you need to edit /etc/pcmcia/config to tell cardmgr which extra modules need to be loaded when your adapter is configured. For example:

     device "aha152x_cs"
       class "scsi" module "scsi/scsi_mod", "scsi/sd_mod", "aha152x_cs"

would say to load the core SCSI module and the top-level disk driver module before loading the regular PCMCIA driver module.

Always turn on SCSI devices before powering up your laptop, or before inserting the adapter card, so that the SCSI bus is properly terminated when the adapter is configured. Also be very careful about ejecting a SCSI adapter. Be sure that all associated SCSI devices are unmounted and closed before ejecting the card. The best way to ensure this is to use either cardctl or cardinfo to request card removal before physically ejecting the card. For now, all SCSI devices should be powered up before plugging in a SCSI adapter, and should stay connected until after you unplug the adapter and/or power down your laptop.

There is a potential complication when using these cards that does not arise with ordinary ISA bus adapters. The SCSI bus carries a “termination power” signal that is necessary for proper operation of ordinary passive SCSI terminators. PCMCIA SCSI adapters do not supply termination power, so if it is required, an external device must supply it. Some external SCSI devices may be configured to supply termination power. Others, such as the Zip Drive and the Syquest EZ-Drive, use active terminators that do not depend on it. In some cases, it may be necessary to use a special terminator block such as the APS SCSI Sentry 2, which has an external power supply. When configuring your SCSI device chain, be aware of whether or not any of your devices require or can provide termination power.

The Adaptec APA-460 SlimSCSI adapter is not supported. This card was originally sold under the Trantor name, and when Adaptec merged with Trantor, they continued to sell the Trantor card with an Adaptec label. The APA-460 is not compatible with any existing Linux driver. I’m not sure how hard it would be to write a driver; I don’t think anyone has been able to obtain the technical information from Adaptec.

The (unsupported) Trantor SlimSCSI can be identified by the following:

     Trantor / Adaptec APA-460 SlimSCSI
     FCC ID: IE8T460
     Shipped with SCSIworks! driver software

The (supported) Adaptec SlimSCSI can be identified by the following:

     Adaptec APA-1460 SlimSCSI
     FCC ID: FGT1460
     P/N: 900100
     Shipped with EZ-SCSI driver software

3.6 How do I use my PCMCIA memory card?

The default memory card startup script will create block and character devices for accessing each partition on a memory card. There are two memory card drivers... pcmem cs, the older driver, works well for simple static RAM cards. The newer driver, memory cs, is mostly for direct access to flash memory cards. Check the man pages for a detailed description of how devices are named. Both block and character devices are created. The block device is used for disk-like access (creating and mounting filesystems, etc). The character device is for ”raw” reads and writes at arbitrary locations.

To use a flash memory card as an ordinary disk-like block device, first create a “flash translation layer” partition on the device with the ftl format command:

     ftl_format -i /dev/mem0c0c

Note that this command accesses the card through the “raw” memory card interface. Once formatted, the card can be accessed as an ordinary block device via the ftl cs driver. For example:

     mke2fs /dev/ftl0
     mount -t ext2 /dev/ftl0 /mnt

With the FTL and “new” memory drivers, the device address passed to ftl.opts and memory.opts consists of two fields: the socket number, and the partition number. Common memory partitions are numbered before attribute memory partitions. Generally, the only interesting partition is partition 0 (the main commonmemory partition, where data is stored). Here is an example of a script that will automatically mount flash memory cards based on which socket they are inserted into:

     case "$ADDRESS" in
     *,0,0)
         # Mount filesystem, but don't update /etc/fstab
         DO_FSTAB="n" ; DO_FSCK="y" ; DO_MOUNT="y"
         FSTYPE="ext2" ; OPTS=""
         MOUNTPT="/ftl0"
         ;;
     *,1,0)
         # Mount filesystem, but don't update /etc/fstab
         DO_FSTAB="n" ; DO_FSCK="y" ; DO_MOUNT="y"
         FSTYPE="ext2" ; OPTS=""
         MOUNTPT="/ftl1"
         ;;
     esac

There are two major formats for flash memory cards: the “flash translation layer” style, and the Microsoft Flash File System. The FTL format is generally more flexible because it allows any ordinary high-level filesystem (ext2, ms-dos, etc) to be used on a flash card as if it were an ordinary disk device. The FFS is a complete new filesystem type. Linux cannot currently handle cards formated with FFS.

3.7 How do I use my ATA/IDE card drive?

ATA/IDE drive support requires a 1.3.72 or higher kernel. The PCMCIA-specific part of the driver is fixed cs. Be sure to use cardctl or cardinfo to shut down an ATA/IDE card before ejecting it, as the driver has not been made “hot-swap-proof”.

The device addresses passed to fixed.opts consist of either two or three fields: the current scheme, the socket number, and an optional partition number. As with SCSI devices, fixed.opts is first called for the entire device. If fixed.opts returns a list of partitions in the PARTS variable, the script will then be called for each partition.

Here is an example fixed.opts file to mount the first partition of any ATA/IDE card on /mnt.

     case "$ADDRESS" in
     *,*)
         PARTS="1"
         ;;
     *,*,1)
         DO_FSTAB="y" ; DO_FSCK="y" ; DO_MOUNT="y"
         FSTYPE="msdos"
         OPTS=""
         MOUNTPT="/mnt"
         ;;
     esac

3.8 How do I tell cardmgr how to identify a new card?

Assuming that your card is supported by an existing driver, all that needs to be done is to add an entry to /etc/pcmcia/config to tell cardmgr how to identify the card, and which driver(s) need to be linked up to this card. Check the man page for pcmcia for more information about the config file format. If you insert an unknown card, cardmgr will normally record some identification information in the system log that can be used to construct the config entry.

Here is an example of how cardmgr will report an unsupported card in /usr/adm/messages.

     cardmgr[460]: unsupported card in socket 1
     cardmgr[460]: version info: "MEGAHERTZ", "XJ2288", "V.34 PCMCIA MODEM"

The corresponding entry in /etc/pcmcia/config would be:

     card "Megahertz XJ2288 V.34 Fax Modem"
       version "MEGAHERTZ", "XJ2288", "V.34 PCMCIA MODEM"
       bind "serial_cs"

You can use “*” to match strings that don’t need to match exactly, like version numbers. When making new config entries, be careful to copy the strings exactly, preserving case and blank spaces. Also be sure that the config entry has the same number of strings as are reported in the log file.

After editing /etc/pcmcia/config, you can signal cardmgr to reload the file with:

     kill -HUP `cat /var/run/cardmgr.pid`

If you do set up an entry for a new card, please send me a copy so that I can include it in the standard config file.

3.9 How do I control which interrupts and ports are used by a device?

In theory, it should not really matter which interrupt is allocated to which device, as long as two devices are not configured to use the same interrupt. In /etc/pcmcia/config.opts you’ll find a place for excluding interrupts that are used by non-PCMCIA devices.

Note that the interrupt used to monitor card status changes is chosen by the low-level socket driver module (i82365 or tcic) before cardmgr parses /etc/pcmcia/config, so it is not affected by changes to this file. To set this interrupt, use the irq mask or cs irq options when the socket driver is loaded, in

/etc/rc.d/rc.pcmcia.

All the client card drivers have a parameter called irq mask for specifying which interrupts they may try to allocate. Each bit of irq mask corresponds to one irq line: bit 0 is irq 0, bit 1 is irq 1, and so on. So, a mask of 0x1200 would correspond to irq 9 and irq 12. To limit a driver to use only one specific interrupt, its irq mask should have only one bit set. These driver options should be set in your /etc/pcmcia/config file. For example:

     device "serial_cs"
       module "serial_cs" opts "irq_mask=0x1100"
       ...

would specify that the serial driver should only use irq 8 or irq 12. Note that Card Services will never allocate an interrupt that is already in use by another device, or an interrupt that is excluded in the config file.

There is no way to directly specify the I/O addresses for a PCMCIA card to use. The /etc/pcmcia/config.opts file allows you to specify ranges of ports available for use by all PCMCIA devices.

After modifying /etc/pcmcia/config, you can restart cardmgr with “kill -HUP”.

3.10 When is it safe to insert or eject a PCMCIA card?

In theory, you can insert and remove PCMCIA cards at any time. However, it is a good idea not to eject a card that is currently being used by an application program. Kernels older than 1.1.77 would often lock up when serial/modem cards were ejected, but this should be fixed now.

3.11 How do I unload PCMCIA drivers?

To unload the entire PCMCIA package, invoke rc.pcmcia with:

     /etc/rc.d/rc.pcmcia stop

This script will take several seconds to run, to give all client drivers time to shut down gracefully. If a PCMCIA device is currently in use, the shutdown will be incomplete, and some kernel modules may not be unloaded. To avoid this, use “cardctl eject” to shut down all sockets before invoking rc.pcmcia. The exit status of the cardctl command will indicate if any sockets could not be shut down.

3.12 How does Card Services deal with suspend/resume?

Card Services can be compiled with support for APM (Advanced Power Management) if you’ve installed this package on your system. APM is incorporated into 1.3.46 and later kernels. It is currently being maintained by Rick Faith (faith@cs.unc.edu), and APM tools can be obtained from ftp.cs.unc.edu in /pub/users/faith/linux. The PCMCIA modules will automatically be configured for APM if a compatible version is detected on your system.

Without resorting to APM, you can do “cardctl suspend” before suspending your laptop, and “cardctl resume” after resuming, to properly shut down and restart your PCMCIA cards. This will not work with a PCMCIA modem that is in use, because the serial driver isn’t able to save and restore the modem operating parameters.

APM seems to be unstable on some systems. If you experience trouble with APM and PCMCIA on your system, try to narrow down the problem to one package or the other before reporting a bug.

3.13 How do I turn off a PCMCIA card without ejecting it?

Use either the cardctl or cardinfo command. “cardctl suspend #” will suspend one socket, and turn off its power. The corresponding resume command will wake up the card in its previous state.

3.14 How can I have separate device configurations for home and work?

This is trivial using PCMCIA “scheme” support. Use two configuration schemes, called “home” and “work”. Here is an example of a network.opts script with scheme-specific settings:

     case "$ADDRESS" in
     work,*,*,*)
         # definitions for network card in work scheme
         ;;
     home,*,*,*|default,*,*,*)
         # definitions for network card in home scheme
         ;;
     esac

The first part of a PCMCIA device address is always the configuration scheme. In this example, the second “case” clause will select for both the “home” and “default” schemes. So, if the scheme is unset for any reason, it will default to the “home” setup.

Now, to choose between the two sets of settings, run either:

     cardctl scheme home

or

     cardctl scheme work

The cardctl command does the equivalent of shutting down all your cards and restarting them. The command can be safely executed whether or not the PCMCIA system is loaded, but the command may fail if you are using other PCMCIA devices at the time (even if their configurations are not explicitly dependant on the scheme setting).

To find out the current PCMCIA scheme setting, run:

     cardctl scheme

4 Problems with specific cards

4.1 Why doesn’t my modem work?

That’s a broad question, but here’s a quick troubleshooting guide.

• Is your card recognized as a modem? Check the system log and make sure that cardmgr identifies the card correctly and starts up the serial cs driver. If it doesn’t, you may need to add a new entry to your /etc/pcmcia/config file so that it will be identified properly. See section 3.8 for details.

• Is the modem configured successfully by serial cs? Again, check the system log and look for messages from the serial cs driver. If you see “register serial() failed”, you may have an I/O port conflict with another device. Another tip-off of a conflict is if the device is reported to be an 8250; most modern PCMCIA modems should be identified as 16550A UART’s. If you think you’re seeing a port conflict, edit /etc/pcmcia/config.opts and exclude the port range that was allocated for the modem.

• Is there an interrupt conflict? If the system log looks good, but the modem just doesn’t seem to work, try using setserial to change the irq to 0, and see if the modem works. This causes the serial driver to use a slower polled mode instead of using interrupts. If this seems to fix the problem, it is likely that some other device in your system is using the interrupt selected by serial cs. You should add a line to /etc/pcmcia/config.opts to exclude this interrupt.

• If the modem seems to work only really, really slowly, this is an almost certain indicator of an interrupt conflict.

• Make sure your problem is really a PCMCIA one. It may help to see if the card works under DOS with the vendor’s drivers. Also, don’t test the card with something complex like SLIP until you are sure you can make simple connections. If simple things work but SLIP does not, your problem is most likely with SLIP, not with PCMCIA.

4.2 Why doesn’t my network adapter card work?

Here’s another quick troubleshooting guide.

• Is your card recognized as an ethernet card? Check the system log and make sure that cardmgr identifies the card correctly and starts up one of the network drivers. If it doesn’t, your card might still be usable if it is compatible with a supported card. This will be most easily done if the card claims to be “NE2000 compatible”.

• Is the card configured properly? If you are using a supported card, and it was recognized by cardmgr, but still doesn’t work, there might be an interrupt or port conflict with another device. Find out what resources the card is using (from the system log), and try excluding these in /etc/pcmcia/config.opts to force the card to use something different.

• If your card seems to be configured properly, but sometimes locks up, particularly under high load, you may need to try changing your socket driver timing parameters. See section 2.3 for more information.

• If you get messages like “network unreachable” when you try to access the network, then you have probably set up /etc/pcmcia/network.opts incorrectly. On the other hand, mis-configured cards will usually fail silently.

• To diagnose problems in /etc/pcmcia/network.opts, start by trying to ping other systems on the same subnet using their IP addresses. Then try to ping your gateway, and then machines on other subnets. Ping machines by name only after trying these simpler tests.

• Make sure your problem is really a PCMCIA one. It may help to see see if the card works under DOS with the vendor’s drivers. Double check your modifications to the /etc/pcmcia/network.opts script. Make sure your drop cable, “T” jack, terminator, etc are working.

Here are some comments about specific cards:

• With IBM CCAE and Socket EA cards, you need to pick the transceiver type (10base2, 10baseT, AUI) when the network device is configured. Make sure that the transceiver type reported in the system log matches your connection.

• The drivers for SMC, Megahertz, Ositech, and 3Com cards should autodetect the attached network type (10base2 or 10baseT). Setting the transceiver type when the driver is loaded serves to define the card’s “first guess”.

• The Farallon EtherWave is actually based on the 3Com 3c589, with a special transceiver. Though the EtherWave uses 10baseT-style connections, its transceiver requires that the 3c589 be configured in 10base2 mode.

• If you have trouble with an IBM CCAE, NE4100, Thomas Conrad, or Kingston adapter, try increasing the memory access time with the mem speed=# option to the pcnet cs module. An example of how to do this is given in the standard config.opts file. Try speeds of up to 1000 (in nanoseconds).

• For the New Media Ethernet adapter, on some systems, it may be necessary to increase the IO port access time with the io speed=# option when the pcmcia core module is loaded. Edit CORE OPTS in

/etc/rc.d/rc.pcmcia to set this option.

• The driver used by the IBM and 3Com token ring adapters seems to behave very badly if the cards are not connected to a ring when they get initialized. Always connect these cards to the net before they are powered up.

4.3 How do I select the transceiver type for a network card?

The transceiver type can be selected in network.opts using the IF PORT setting. This can either be a numeric value as in previous PCMCIA releases, or a keyword identifying the transceiver type. All the network drivers default to either autodetect the interface if possible, or 10baseT otherwise. The ifport command can be used to check or set the current transceiver type. For example:

# ifport eth0 10base2
#
# ifport eth0
eth0    2 (10base2)

Current releases of the 3c589 driver attempt to autodetect the network connection, but this doesn’t seem to be completely functional yet. For autodetection to work, the network cable should be connected to the card when the card is configured. Alternatively, once the network is connected, you can force the driver to check the connection with:

ifconfig eth0 down up

4.4 How do I add support for an NE2000-compatible ethernet card?

First, see if the card is already recognized by cardmgr. Some cards not listed in SUPPORTED.CARDS are actually OEM versions of cards that are supported. If you find a card like this, let me know so I can add it to the list.

If your card is not recognized, follow the instructions in section 3.8 to create a config entry for your card, but bind the card to the memory card driver, pcmem cs for now. Restart cardmgr to use the new updated config file.

You will need to know your card’s hardware ethernet address. This address is a series of six two-digit hex numbers, often printed on the card itself. If it is not printed on the card, you may be able to use a DOS driver to display the address. In any case, once you know it, run:

     dd if=/dev/pcmem0a count=20 | od -Ax -t x1

and search the output for your address. Record the hex offset of the first byte of the address. Now, edit modules/pcnet cs.c and find the hw info structure. You’ll need to create a new entry for your card. The first field is a descriptive name. The next field is the offset multiplied by two. The next three fields are the first three bytes of the hardware address. The final field contains some flags for specific card features; to start, try setting it to 0.

After editing pcnet cs.c, compile and install the new module. Edit /etc/pcmcia/config again, and change the card binding from pcmem cs to pcnet cs. Follow the instructions for reloading the config file, and you should be all set. Please send me copies of your new hw info and config entries.

If you can’t find your card’s hardware address in the hex dump, as a method of last resort, it is possible to “hardwire” the address when the pcnet cs module is initialized. Edit /etc/pcmcia/config and add a hw addr= option, like so:

     module "pcnet_cs" opts "hw_addr=0x00,0x80,0xc8,0x01,0x02,0x03"

Substitute your own card’s hardware address in the appropriate spot, of course.

4.5 Multifunction cards such as the 3Com 3c562

Starting with the 1.3.73 Linux kernel, a single interrupt can be shared by several drivers, such as the serial driver and an ethernet driver. When using a multifunction card under a newer kernel, all card functions can be used simultaneously.

Earlier kernels do not support interrupt sharing between different device drivers, so it is not possible for the PCMCIA drivers to configure this card for simultaneous ethernet and modem access. The ethernet and serial drivers are both loaded automatically. However, the ethernet driver “owns” the card interrupt by default. To use the modem, you can unload the ethernet driver and reconfigure the serial port by doing something like:

     ifconfig eth0 down

     rmmod 3c589_cs
     setserial /dev/modem autoconfig auto_irq
     setserial /dev/modem

The second setserial should verify that the port has been configured to use the interrupt previously used by the ethernet driver.

4.6 How do I use my PCMCIA floppy interface?

The PCMCIA floppy interface used in the Compaq Aero and a few other laptops is not yet supported by this package. The snag in supporting the Aero floppy is that the Aero seems to use a customized PCMCIA controller to support DMA to the floppy. Without knowing exactly how this is done, there isn’t any way to implement support under Linux.

If the floppy adapter card is present when an Aero is booted, the Aero BIOS will configure the card, and Linux will identify it as a normal floppy drive. When the Linux PCMCIA drivers are loaded, they will notice that the card is already configured and attached to a Linux driver, and this socket will be left alone. So, the drive can be used if it is present at boot time, but the card is not hot swappable.

4.7 What’s up with support for Xircom cards?

Xircom does not share technical information about its cards without a non-disclosure agreement. This means that it is not really possible to develop freely distributable drivers for Xircom cards without doing legally dubious things like reverse engineering DOS drivers.

Recently, Xircom seems to have had a change of heart about their non-disclosure policy, and it now seems that this information will be more readily available in the future. However, please do not pester me with questions about Xircom cards: drivers will be announced when they become available, and I will not try to guess when that will be.

The Xircom CreditCard Ethernet+Modem II card can be used as a modem under Linux, with no special configuration.

5 Debugging tips and programming information

5.1 How can I submit a helpful bug report?

Here are some things that should be included in all bug reports:

• Your system type, and the output of the probe command

• What PCMCIA cards you are using

• Your Linux kernel version, and PCMCIA version

• The contents of the config.out file • Any changes you’ve made to the startup files in /etc/pcmcia, or to rc.pcmcia

• All PCMCIA-related stuff in your system log file

Before submitting a bug report, please check to make sure that you are using an up-to-date copy of the driver package. While it is somewhat gratifying to read bug reports for things I’ve already fixed, it isn’t a particularly constructive use of my time.

If your problem involves a kernel fault, the register dump from the fault is only useful if you can track down the fault address, EIP. If it is in the main kernel, look up the address in System.map to identify the function at fault. If the fault is in a loadable module, it is a bit harder to trace. With the current module tools, “ksyms -m” will report the base address of each loadable module. Pick the module that contains the EIP address, and subtract its base address from EIP to get an offset inside that module. Then, run gdb on that module, and look up the offset with the list command. This will only work if you’ve compiled that module with -g to include debugging information.

Send bug reports to dhinds@allegro.stanford.edu. I prefer to handle bug reports by email – please avoid calling me at home or at work. You can also submit bug reports via the WWW: see

http://hyper.stanford.edu/~dhinds/pcmcia/pcmcia.html for details.

5.2 Low level PCMCIA debugging aids

The PCMCIA modules contain a lot of conditionally-compiled debugging code. Most of this code is under control of the PCMCIA DEBUG preprocessor define. If this is undefined, debugging code will not be compiled. If set to 0, the code is compiled but inactive. Larger numbers specify increasing levels of verbosity. Each module built with PCMCIA DEBUG defined will have an integer parameter, pc debug, that controls the verbosity of its output. This can be adjusted when the module is loaded, so output can be controlled on a per-module basis without recompiling.

There are a few debugging tools in the debug tools/ subdirectory of the PCMCIA distribution. The dump tcic and dump i365 utilities generate complete register dumps of the PCMCIA controllers, and decode a lot of the register information. They are most useful if you have access to a datasheet for the corresponding controller chip. The dump tuples utility lists a card’s CIS (Card Information Structure), and decodes some of the important bits. And the dump cisreg utility displays a card’s local configuration registers.

The pcmem cs memory card driver is also sometimes useful for debugging. It can be bound to any PCMCIA card, and does not interfere with other drivers. It can be used to directly access any card’s attribute memory or common memory.

5.3 How do I write a Card Services driver for card X?

The Linux PCMCIA Programmer’s Guide is the best documentation for the Linux PCMCIA interface. The latest version is always available from cb-iris.stanford.edu in /pub/pcmcia/doc.

For devices that are close relatives of normal ISA devices, you’ll probably be able to use parts of existing Linux drivers. In some cases, the biggest stumbling block will be modifying an existing driver so that it can handle adding and removing devices after boot time. Of the current drivers, the memory card driver is the only “self-contained” driver that does not depend on other parts of the Linux kernel to do most of the dirty work.

I’ve written a skeleton driver with lots of comments that explains a lot of how a driver communicates with Card Services; you’ll find this in the PCMCIA source distribution in modules/skeleton.c.